# A Note on Restricted Selection Set from Random Interval

### v1

@@ 2.7. Event probabilities: unrestricted and mean-restricted ranges
Corollary 2.15 (Identification formula for $\mathbf{P}(y^{*}\in A)$).
Let $y^{\star}:=y_{\lambda^{\star}}$ be an optimizer from Theorem˜2.14, and set $t^{\star}:=1/\lambda^{\star}$ if $\lambda^{\star}>0$ and $s^{\star}:=1/(-\lambda^{\star})$ if $\lambda^{\star}<0$. Then
[…]
and the thresholds are pinned down by the mean-calibration equations
$\mathbf{E}\!\big[y_{U}-\min\{\Delta^{+},t^{\star}\}\big]=\kappa\quad(\lambda^{\star}>0),\qquad\mathbf{E}\!\big[y_{L}+\min\{\Delta^{-},s^{\star}\}\big]=\kappa\quad(\lambda^{\star}<0).$ (2.19)
@@
For $\lambda^{\star}>0$, the optimizer in Theorem 2.14 has the form
$y^{\star}(\omega)=\begin{cases}y_{U}(\omega),&Y(\omega)\cap A=\emptyset,\\[2.0pt] y_{U}(\omega)-\min\{\Delta^{+}(\omega),t^{\star}\},&Y(\omega)\cap A\neq\emptyset,\end{cases}$
with $t^{\star}=1/\lambda^{\star}$. On $\{Y\cap A\neq\emptyset\}$, when $\Delta^{+}(\omega)<t^{\star}$ we have $y^{\star}(\omega)=a^{+}(\omega)\in A$; when $\Delta^{+}(\omega)>t^{\star}$ we have $y^{\star}(\omega)=y_{U}(\omega)\notin A$ (except possibly on the boundary of $A$). On the level set $\{\Delta^{+}=t^{\star}\}$ both choices yield the same value of the Lagrangian, so tie-breaking can be used to match the mean constraint while keeping the probability as in (2.18). The mean calibration (2.19) is exactly the equation $\mathbb{E}[y^{\star}]=\kappa$ restricted to the case $\lambda^{\star}>0$. The case $\lambda^{\star}<0$ is analogous, using the expression for $y_{\lambda}$ in terms of $y_{L}$ and $\Delta^{-}$. When $\lambda^{\star}=0$, the mean constraint is nonbinding in the Lagrangian and the optimal strategy is simply to choose a selection that hits $A$ whenever possible, giving probability $\mathbf{P}(Y\cap A\neq\emptyset)$. ∎
@@
$\Delta^{+}(\omega)=\begin{cases}\big(y_{U}(\omega)-b\big)_{+},&y_{L}(\omega)\leq b,\\ +\infty,&y_{L}(\omega)>b,\end{cases}\qquad\Delta^{-}(\omega)=\begin{cases}\big(a-y_{L}(\omega)\big)_{+},&y_{U}(\omega)\geq a,\\ +\infty,&y_{U}(\omega)<a,\end{cases}$
so the threshold rule (2.17) switches at a deterministic distance from the right edge $b$ in the high-mean regime and from the left edge $a$ in the low-mean regime, with $t^{\star}$ or $s^{\star}$ determined by (2.19).
@@ 2.8. Further extensions: higher moments and additional quantiles
over selections $y$ of $Y$. By an Artstein-type interchange theorem for random closed sets (see [Artstein1974, Molchanov2005]),
@@
cf. [Artstein1974, Molchanov2005]. Standard convex duality arguments (see, e.g., [Rockafellar1970]) yield
@@
For $r=1$ and $\mu_{1}=\kappa$ we have $\operatorname{\mathbf{Sel}}_{1}(Y\mid\mu_{1})=\operatorname{\mathbf{Sel}}(Y\mid\kappa)$ and $\underline{m}(1,\mu_{1})=\overline{m}(1,\mu_{1})=\kappa$, recovering the trivial point identification of the mean. For $r\neq 1$, the interval $\underline{m}(r,\mu_{r}),\overline{m}(r,\mu_{r})]$ describes the shrinkage of the benchmark range $\mathbb{E}y_{L},\mathbb{E}y_{U}]$ induced by the $r$th-moment restriction.
@@
We now use this to show that $\Theta_{E}(\alpha,q_{\alpha})$ is convex. The proof formally works on a product extension that introduces independent randomization, but because our identification analysis only depends on the joint law of $(Y,y)$, the attainable range of $\mathbb{E}[y]$ is unaffected by such an enlargement.
@@
Definition 2.26 (Quantile-identified mean interval).
@@
will be called the $(\alpha,q_{\alpha})$-quantile identified mean interval.

### v2

@@ 2.7. Event probabilities: unrestricted and mean-restricted ranges
Corollary 2.15 (Extremal probability formula).
Let $\lambda^{\star}\in\mathbb{R}$ and let
$y^{\star}\in\operatorname{\mathbf{Sel}}(Y_{\lambda^{\star}})$
be any optimizer provided by Theorem 2.14, so that
$\mathbb{E}[y^{\star}]=\kappa,\qquad\mathbf{P}(y^{\star}\in A)=U(\kappa).$
Set
$t^{\star}:=1/\lambda^{\star}\qquad\text{if }\lambda^{\star}>0,\qquad s^{\star}:=1/(-\lambda^{\star})\qquad\text{if }\lambda^{\star}<0.$
Then
$\mathbf{P}(y^{\star}\in A)=\begin{cases}\mathbf{P}\!\big(\Delta^{+}<t^{\star}\big)+\eta^{+}\,\mathbf{P}\!\big(\Delta^{+}=t^{\star}\big),&\lambda^{\star}>0,\\[4.0pt] \mathbf{P}\!\big(\Delta^{-}<s^{\star}\big)+\eta^{-}\,\mathbf{P}\!\big(\Delta^{-}=s^{\star}\big),&\lambda^{\star}<0,\\[4.0pt] \mathbf{P}\big(Y\cap A\neq\varnothing\big),&\lambda^{\star}=0,\end{cases}$ (2.22)
for some $\eta^{+},\eta^{-}\in[0,1]$ determined by the tie-breaking on the boundary sets. Equivalently, the optimal probability is obtained by full inclusion of the strict-threshold region and possible randomization on the tie set.
Moreover, the threshold parameters are pinned down by the mean-calibration conditions
$\kappa\in\Big[\mathbb{E}\!\big[y_{U}-\min\{\Delta^{+},t^{\star}\}\big],\,\mathbb{E}\!\big[y_{U}-\min\{\Delta^{+},t^{\star}\}\big]+\frac{1}{\lambda^{\star}}\mathbf{P}(\Delta^{+}=t^{\star})\Big]\qquad(\lambda^{\star}>0),$ (2.23)
and
$\kappa\in\Big[\mathbb{E}\!\big[y_{L}+\min\{\Delta^{-},s^{\star}\}\big],\,\mathbb{E}\!\big[y_{L}+\min\{\Delta^{-},s^{\star}\}\big]+\frac{1}{-\lambda^{\star}}\mathbf{P}(\Delta^{-}=s^{\star})\Big]\qquad(\lambda^{\star}<0).$ (2.24)
In particular, if
$\mathbf{P}(\Delta^{+}=t^{\star})=0\quad\text{or}\quad\mathbf{P}(\Delta^{-}=s^{\star})=0,$
then no tie-breaking is needed and the calibration conditions reduce to the equalities
$\mathbf{E}\!\big[y_{U}-\min\{\Delta^{+},t^{\star}\}\big]=\kappa\qquad(\lambda^{\star}>0),$
and
$\mathbf{E}\!\big[y_{L}+\min\{\Delta^{-},s^{\star}\}\big]=\kappa\qquad(\lambda^{\star}<0).$
@@
Suppose first that $\lambda^{\star}>0$. By Theorem 2.14, the tie set is
$T_{\lambda^{\star}}^{+}=\{\omega:Y(\omega)\cap A\neq\emptyset,\ \lambda^{\star}\Delta^{+}(\omega)=1\}=\{\omega:\Delta^{+}(\omega)=t^{\star}\},$
and outside this set every maximizer is uniquely given by
$y^{\star}(\omega)=y_{U}(\omega)-\min\{\Delta^{+}(\omega),t^{\star}\}.$
Hence:
$\Delta^{+}(\omega)<t^{\star}\Longrightarrow y^{\star}(\omega)=a^{+}(\omega)\in A,$
$\Delta^{+}(\omega)>t^{\star}\Longrightarrow y^{\star}(\omega)=y_{U}(\omega)\notin A,$
while on $\{\Delta^{+}=t^{\star}\}$ both $a^{+}(\omega)\in A$ and $y_{U}(\omega)\notin A$ are pointwise maximizers. Therefore
$\mathbf{P}(y^{\star}\in A)=\mathbf{P}(\Delta^{+}<t^{\star})+\eta^{+}\mathbf{P}(\Delta^{+}=t^{\star})$
for some $\eta^{+}\in[0,1]$, depending on the measurable tie-breaking on the boundary set.
For the mean, the minimal choice among maximizers is
$y_{\star}^{-}(\omega):=y_{U}(\omega)-\min\{\Delta^{+}(\omega),t^{\star}\},$
which selects $a^{+}$ on the tie set, while the maximal choice selects $y_{U}$ on the tie set. Since on $\{\Delta^{+}=t^{\star}\}$,
$y_{U}(\omega)-a^{+}(\omega)=\Delta^{+}(\omega)=t^{\star}=\frac{1}{\lambda^{\star}},$
the set of attainable means across measurable tie-breakings is exactly
$\Big[\mathbb{E}\!\big[y_{U}-\min\{\Delta^{+},t^{\star}\}\big],\,\mathbb{E}\!\big[y_{U}-\min\{\Delta^{+},t^{\star}\}\big]+\frac{1}{\lambda^{\star}}\mathbf{P}(\Delta^{+}=t^{\star})\Big].$
Since $y^{\star}$ is chosen so that $\mathbb{E}[y^{\star}]=\kappa$, this proves (2.23).
The case $\lambda^{\star}<0$ is analogous. Now the tie set is
$T_{\lambda^{\star}}^{-}=\{\omega:Y(\omega)\cap A\neq\emptyset,\ (-\lambda^{\star})\Delta^{-}(\omega)=1\}=\{\omega:\Delta^{-}=s^{\star}\},$
and outside this set every maximizer is uniquely given by
$y^{\star}(\omega)=y_{L}(\omega)+\min\{\Delta^{-}(\omega),s^{\star}\}.$
Thus
$\mathbf{P}(y^{\star}\in A)=\mathbf{P}(\Delta^{-}<s^{\star})+\eta^{-}\mathbf{P}(\Delta^{-}=s^{\star})$
for some $\eta^{-}\in[0,1]$. Moreover, on $\{\Delta^{-}=s^{\star}\}$,
$a^{-}(\omega)-y_{L}(\omega)=\Delta^{-}(\omega)=s^{\star}=\frac{1}{-\lambda^{\star}},$
so the attainable means across measurable tie-breakings form exactly the interval
$\Big[\mathbb{E}\!\big[y_{L}+\min\{\Delta^{-},s^{\star}\}\big],\,\mathbb{E}\!\big[y_{L}+\min\{\Delta^{-},s^{\star}\}\big]+\frac{1}{-\lambda^{\star}}\mathbf{P}(\Delta^{-}=s^{\star})\Big],$
which gives (2.24).
Finally, if $\lambda^{\star}=0$, then by Theorem 2.14,
$Y_{0}(\omega)=\begin{cases}Y(\omega)\cap A,&Y(\omega)\cap A\neq\emptyset,\\ Y(\omega),&Y(\omega)\cap A=\emptyset,\end{cases}$
so every optimizer chooses a point in $A$ whenever possible. Hence
$\mathbf{P}(y^{\star}\in A)=\mathbf{P}(Y\cap A\neq\emptyset).$
This completes the proof. ∎
@@
$\Delta^{+}(\omega)=\begin{cases}\bigl(y_{U}(\omega)-b\bigr)_{+},&y_{L}(\omega)\leq b,\\ +\infty,&y_{L}(\omega)>b,\end{cases}\qquad\Delta^{-}(\omega)=\begin{cases}\bigl(a-y_{L}(\omega)\bigr)_{+},&y_{U}(\omega)\geq a,\\ +\infty,&y_{U}(\omega)<a.\end{cases}$
Hence, in the high-mean regime $\lambda^{\star}>0$, the optimizer is described by (2.18): it switches from $y_{U}$ to the rightmost point of $Y\cap A$ once the gap to the right edge $b$ is small enough. In the low-mean regime $\lambda^{\star}<0$, the optimizer is described by (2.19): it switches from $y_{L}$ to the leftmost point of $Y\cap A$ once the gap to the left edge $a$ is small enough. The corresponding threshold $t^{\star}=1/\lambda^{\star}$ or $s^{\star}=1/(-\lambda^{\star})$ is then pinned down by the mean-calibration equations (2.23)–(2.24).
@@ 2.8. Further extensions: higher moments and additional quantiles
over selections $y$ of $Y$. By an Artstein-type interchange theorem for random closed sets (see [1, 7]),
@@
cf. [1, 7]. Standard convex duality arguments (see, e.g., [8]) yield
@@
For $r=1$ and $\mu_{1}=\kappa$ we have $\operatorname{\mathbf{Sel}}_{1}(Y\mid\mu_{1})=\operatorname{\mathbf{Sel}}(Y\mid\kappa)$ and $\underline{m}(1,\mu_{1})=\overline{m}(1,\mu_{1})=\kappa$, recovering the trivial point of the mean. For $r\neq 1$, the interval $\underline{m}(r,\mu_{r}),\overline{m}(r,\mu_{r})]$ describes the shrinkage of the benchmark range $\mathbb{E}y_{L},\mathbb{E}y_{U}]$ induced by the $r$th-moment restriction.
@@
We now use this to show that $\Theta_{E}(\alpha,q_{\alpha})$ is convex. The proof formally works on a product extension that introduces independent randomization, but because our analysis only depends on the joint law of $(Y,y)$, the attainable range of $\mathbb{E}[y]$ is unaffected by such an enlargement.
@@
Definition 2.26 (quantile-restricted mean interval).
@@
will be called the $(\alpha,q_{\alpha})$-quantile-restricted mean interval.